Find: 1/7 of 6/5 Find: 1/7 of 6/5 Find: 1/7 of 6/5 Find: 1/7 of 6/5

 Find:
1/7 of 6/5
 Find:
1/7 of 6/5
 Find:
1/7 of 6/5
 Find:
1/7 of 6/5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (1/7) × (6/5)
By the rule Multiplication of fraction,
Product of fraction = (product of numerator)/ (product of denominator)
Then,
= (1/7) × (6/5)
= (1 × 6)/ (7 × 5)
= (6/35)

Thanks